Eight uncharged capacitors with equal capacitances are combined in parallel. The combination is connected to a 6.85V battery , which charges the capacitors. The charging process involves 0.000381 C of charge moving through the battery. Find the capacitance of each capacitor

Physicist Worries xLakehead University-PHYS 1xN Necis Lyndsay Williams Jump to.. ioi i33-Winter ї9 . MACKAY > Actiities and Due Dates > MW: Electric Potential Resources QGive up? Feedback Resame nt Score: 1345/2500 Attempt 1 İof25 parallel. The combination is connected to a 6.85 V Eight uncharged capacitors with equal capacitances are combined i battery, which charges the capacitors. The charging process involves 0.000381 C of charge moving through the hattery. Find the capacitance of each capacitor capacitance: 445 x104 5 7 8
Answer #1

Here , let each capacitor is C

for net capacitance , Cnet = 8 * C

for the net charge

charge = 8 * C * Voltage

0.000381 = 8 * C * 6.85

solving for C

C = 6.95 *10^-6 C

the capacitance of each is 6.95 *10-6 C